# A Survey on Multilingual Mental Disorders Detection from Social Media Data

### v1

@@ Abstract
The increasing prevalence of mental health disorders globally highlights the urgent need for effective digital screening methods that can be used in multilingual contexts. Most existing studies, however, focus on English data, overlooking critical mental health signals that may be present in non-English texts. To address this important gap, we present the first survey on the detection of mental health disorders using multilingual social media data. We investigate the cultural nuances that influence online language patterns and self-disclosure behaviors, and how these factors can impact the performance of NLP tools. Additionally, we provide a comprehensive list of multilingual data collections that can be used for developing NLP models for mental health screening. Our findings can inform the design of effective multilingual mental health screening tools that can meet the needs of diverse populations, ultimately improving mental health outcomes on a global scale.
@@ A Survey on Multilingual Mental Disorders Detection
Ana-Maria Bucur1,2,3, Marcos Zampieri4, Tharindu Ranasinghe5, Fabio Crestani3 1Interdisciplinary School of Doctoral Studies, University of Bucharest, Romania 2PRHLT Research Center, Universitat Politècnica de València, Spain 3Università della Svizzera italiana, Switzerland 4George Mason University, USA 5Aston University, Birmingham, UK
@@ 1 Introduction
It is estimated that nearly half of the population will develop at least one mental disorder by the age of 75 McGrath et al. (2023). Unfortunately, many people do not seek psychiatric help for mental health issues due to stigma, which manifests itself differently between cultures and is influenced by different cultural norms, religious beliefs and social attitudes Ahad et al. (2023). Due to the stigma associated with mental health and the limited access to professional care around the world, the World Health Organization (WHO) advocates for improved delivery of mental health services, including digital technologies to deliver remote care. [^1] https://www.who.int/news/item/17-06-2022-who-highlights-urgent-need-to-transform-mental-health-and-mental-health-care There is a pressing need for the integration of remote screening tools and the delivery of culturally adapted digital interventions Bond et al. (2023). Remote screening relies on processing language patterns associated with mental disorders, which can be identified from short essay writing Rude et al. (2004), text messages Nobles et al. (2018), or social media Eichstaedt et al. (2018).
@@
The first well-known study on the detection of mental disorders using social media was conducted by De Choudhury et al. (2013). Multiple other studies have shown that the language used on Facebook can predict future depression diagnoses found in medical records, indicating that social media data could serve as a valuable complement to depression screening Eichstaedt et al. (2018). The current methods used for social media screening focus mainly on English data Skaik and Inkpen (2020); Harrigian et al. (2021). Additionally, there have been multiple workshops and shared tasks addressing NLP applications to mental health primarily on English data such as eRisk Parapar et al. (2024), CLPsych Chim et al. (2024) and LT-EDI Kayalvizhi et al. (2023).
@@
There are important limitations in current NLP models when processing multilingual mental health-related data. Various studies analyzing English data from social media have shown that there are cultural differences in online language markers of mental disorders De Choudhury et al. (2017); Aguirre and Dredze (2021); Rai et al. (2024) and that the NLP models used for detection do not generalize on data from non-Western cultures Aguirre et al. (2021); Abdelkadir et al. (2024). Even one of the best predictors of depression in language, the use of the first person pronoun "I" Rude et al. (2004), for example, has different degrees of association with the severity of depression across different demographic groups Rai et al. (2024). This suggests that markers of mental disorders in social media language are not universal. One reason for this variation is that self-disclosure rates differ between cultures; collectivist cultures tend to have lower self-disclosure rates than individualist cultures in online settings Tokunaga (2009). Furthermore, non-native English speakers tend to use their native language for more intimate self-disclosures on social media, with higher rates of negative disclosure compared to posts in English Tang et al. (2011). This could have substantial implications for English-based social media screening tools, as they can overlook important signals of mental health disorders that are present in posts that are not written in English.
@@
Recently, there have been efforts to develop detection models that focus on languages other than English, such as Portuguese Santos et al. (2024), German Zanwar et al. (2023), Arabic Almouzini et al. (2019), and Chinese Zhu et al. (2024). There have also been shared tasks specifically designed to address these issues, such as MentalRiskES Mármol-Romero et al. (2023), which focuses on the early detection of depression, suicide, and eating disorders in Spanish. To further contribute to these important efforts, we present the first survey on mental disorders detection from multilingual social media data. This survey aims to promote the development of multilingual NLP models that take into account cross-cultural and cross-language differences in online language.
@@
    We investigate cross-cultural and cross-language differences in the manifestations of mental disorders in social media.
@@
    We provide a comprehensive list of multilingual mental health datasets that capture linguistic diversity and can be used for developing multilingual NLP models. [^2] We make the list available and we will continuously update it: https://github.com/bucuram/multilingual-mental-health-datasets-nlp
@@
    We identify and describe several research gaps and future directions in the detection of multilingual mental disorders using online data.
@@ 2 Related Surveys
In this section, we analyze related surveys on the analysis of mental disorders from social media data. Calvo et al. (2017) is considered one of the first comprehensive surveys, presenting the datasets and NLP techniques used for mental health status detection and intervention. The survey explores research on various mental health conditions and states, including depression, mood disorders, psychological distress, and suicidal ideation, specifically in non-clinical texts such as user-generated content from social media and online forums. Similarly, recent surveys from Skaik and Inkpen (2020); Harrigian et al. (2021); Ríssola et al. (2021); Zhang et al. (2022); Garg (2023); Bucur et al. (2025) present the datasets, features, and models used to detect mental disorders from online content, focusing mainly on English language data. In addition to these surveys, Chancellor and De Choudhury (2020) provides a critical review of the study design and methods used to predict mental health status, along with recommendations to improve research in this field. Dhelim et al. (2023); Bucur et al. (2025) focus on studies that were published during the COVID-19 pandemic. It focuses on general mental well-being, loneliness, anxiety, stress, PTSD, depression, suicide, and other mental disorders.
Our paper fills an important gap in the literature by offering the first comprehensive survey of research on detecting mental disorders in languages other than English. The most related survey to ours is the one by Garg (2024), which focuses exclusively on low-resource languages. Our survey, however, has a broader scope as it discusses work on many languages irrespective of their resourcefulness.
@@ 3 Mental Disorders Detection Tasks Overview
In this section, we discuss the most common tasks related to predicting mental health disorders. When available, we include references to studies that focus on languages other than English. The prediction of mental health issues through social media is typically approached as a supervised classification task (Figure 1). The most common focus is on the binary classification of mental disorders. In this process, a collection of social media posts is used to train an NLP model, which then predicts a binary label that indicates the presence or absence of a mental disorder. Binary classification can be performed at the post-level, which is often used to predict suicidal ideation Huang et al. (2019) and depression Uddin et al. (2019). However, relying solely on a single post for decision making can lead to inaccurate predictions. Therefore, predictions can be made at the user level to detect conditions like depression Hiraga (2017), anxiety Zarate et al. (2023), bipolar disorder Sekulić et al. (2018), etc. Binary classification at the user level can also be modeled as an early risk prediction task, which aims to accurately label users as soon as possible, allowing the model to make a prediction or wait for more data before deciding Losada and Crestani (2016); Parapar et al. (2021).
Another important task is severity prediction, which can be modeled either as an ordinal regression / classification task or as a multiclass classification task. It is used primarily to predict the severity of depression Naseem et al. (2022); Kabir et al. (2023); Sampath and Durairaj (2022) or the risk of suicide attempts Benjachairat et al. (2024). Social media posts can be modeled longitudinally to detect moments of change in the mental health status of individuals. These shifts or escalations in mood can be used as a warning signal for potential suicidal behavior Tsakalidis et al. (2022b).
There are tasks designed to improve the explainability of the field, such as symptom prediction for mental disorders Liu et al. (2023); Yadav et al. (2020). Another step toward improving the explainability of model predictions is highlighting evidence for mental disorders Chim et al. (2024); Varadarajan et al. (2024). Mental health indicators from the social media timeline of an individual can be used to fill in validated questionnaires, with the goal of estimating symptoms of mental disorders that are usually assessed through survey-based methods such as the Beck’s Depression Inventory-II (BDI-II) [^3] https://naviauxlab.ucsd.edu/wp-content/uploads/2020/09/BDI21.pdf for depression assessment Parapar et al. (2021) or the Eating Disorder Examination Questionnaire (EDE-Q) [^4] https://www.corc.uk.net/media/1273/ede-q_quesionnaire.pdf for eating disorders Parapar et al. (2024).
Finally, mental health monitoring aggregated results from detection systems can be used to estimate the prevalence of mental disorders within a population. This approach was used during the COVID-19 pandemic to assess mental health burden, with results comparable to traditional survey-based methods Cohrdes et al. (2021).
4 Shared Tasks
Shared tasks have encouraged interdisciplinary collaborations between psychologists and computer scientists, resulting in systems that help detect mental disorders through social media analysis. These shared tasks have provided benchmark datasets that the research community continues to use, even beyond the official competitions.
The Early Detection of Mental Disorders Risk in Spanish (MentalRiskES) is the only shared task focused on detecting mental disorders in languages other than English. MentalRiskES includes tasks such as the detection of depression, anxiety, eating disorders, and suicidal risk in the Spanish language Mármol-Romero et al. (2023).
@@ 4 Shared Tasks
Sinhala Rare Rathnayake and Arachchige (2021), EmoMent Atapattu et al. (2022), Herath and Wijayasiriwardhane (2024)
Table 1: Available multilingual datasets for detecting mental disorders.
Other shared tasks are focused only on social media data in English. The Early Risk Prediction on the Internet Lab (eRisk) is an annual competition focusing mainly on the early detection of mental disorders, including depression, self-harm, pathological gambling, and eating disorders Parapar et al. (2024). The Workshop on Computational Linguistics and Clinical Psychology (CLPsych) includes various tasks, such as detecting depression and PTSD Coppersmith et al. (2015), labeling crisis posts Milne et al. (2016), and identifying moments of change Tsakalidis et al. (2022a) The Workshop on Language Technology for Equality, Diversity, and Inclusion (LT-EDI) organized tasks for predicting the severity of depression Kayalvizhi et al. (2023).
5 Methodology
To identify datasets for modeling the manifestations of mental disorders in languages other than English, we conducted a systematic search on major publication databases, including ACL Anthology, ACM Digital Library, IEEE Xplore, Springer Nature Link, ScienceDirect, and Google Scholar. Initially, 405 studies were identified through database searches. After screening the abstracts, 215 papers were excluded because they did not mention the language of the data, or mention that the data is in English. Thus, following a review of the main body of the papers, the number of eligible studies was narrowed down to 108, which represents the final count of papers presenting datasets. Papers that did not present new data collections in languages other than English were excluded during the screening process. The PRISMA flow diagram for the survey is presented in Figure 3 in the Appendix.
@@ 6 Multilingual Datasets
The languages most frequently represented in the data collections are three high-resource languages: Chinese, Arabic, and Spanish. Although approximately half of the datasets were published in unranked venues, leading to low visibility for the research, the other half were published in high-ranking journals and conferences (Figure 4 in Appendix A.
@@ 6.1 Data Sources
Most of the datasets in English are sourced from Twitter [^5] All the datasets were collected before Twitter changed its name to X, so we refer to it as ‘Twitter’ in this paper. and Reddit Harrigian et al. (2021). Most non-English datasets in this section were also primarily collected from Twitter. However, Reddit was not as widely used for these data collections in non-English contexts. People use social media platforms differently. Twitter provides community and safety, helping raise awareness and combat stigma around mental health Berry et al. (2017). In contrast, Reddit allows for greater anonymity with “throwaway” accounts, encouraging users to openly share their experiences in detailed posts on specific subreddits De Choudhury and De (2014). This longer format supports post-level mental health analysis Chowdhury et al. (2024), while Twitter’s shorter posts favor user-level insights, requiring longitudinal data to identify patterns Tumaliuan et al. (2024). The data presented in this survey come from various populations and regions, and some of the sources are platforms that are exclusive to specific countries, such as Sina Weibo [^6] https://weibo.com used in China, VKontakte [^7] https://vk.com/ used in Russia, Pantip [^8] https://pantip.com/ in Thailand, or Everytime [^9] https://everytime.kr/ in Korea.
@@ 6.2 Languages
Table 1 presents all the datasets with multilingual data. A more detailed version of the table can be found in Appendix A, Table LABEL:tab:datasets. For classifying resource types, we used the framework proposed by Joshi et al. (2020). Figure 4 illustrates that most of the languages used in the data collections belong to some of the largest language families by number of speakers, specifically the Indo-European, Sino-Tibetan and Afro-Asiatic language families. The languages most frequently represented in the data collections are high-resource languages: Chinese appears in 25 data collections, Arabic is found in 11 datasets, and Spanish is included in 10 datasets. Even if most of the languages covered in the data are from high-, mid to high- and mid-resourced languages, we also have some languages with fewer resources, such as Cantonese and Norwegian. The Cantonese data collection was gathered by Gao et al. (2019) from Youtube comments and annotated for the risk of suicide. The Norwegian datasets related to depression were collected from a public online forum in Norway Uddin et al. (2022); Uddin (2022). Sinhala language, which was classified as rare by Joshi et al. (2020) is represented in three research papers. One of the papers contains Facebook data annotated for suicide ideation Herath and Wijayasiriwardhane (2024), while another contains depression-related data from Twitter and Facebook Rathnayake and Arachchige (2021). The third dataset contains data from Facebook, with more fine-grained labeled data on the presence of mental illness, anxiety, suicidal ideation, emotions, psychosomatic symptoms, and other manifestations Atapattu et al. (2022).

### v2

@@ Abstract
The increasing prevalence of mental disorders globally highlights the urgent need for effective digital screening methods that can be used in multilingual contexts. Most existing studies, however, focus on English data, overlooking critical mental health signals that may be present in non-English texts. To address this gap, we present a survey of the detection of mental disorders using social media data beyond the English language. We compile a comprehensive list of 108 datasets spanning 25 languages that can be used for developing NLP models for mental health screening. In addition, we discuss the cultural nuances that influence online language patterns and self-disclosure behaviors, and how these factors can impact the performance of NLP tools. Our survey highlights major challenges, including the scarcity of resources for low- and mid-resource languages and the dominance of depression-focused data over other disorders. By identifying these gaps, we advocate for interdisciplinary collaborations and the development of multilingual benchmarks to enhance mental health screening worldwide.
@@ A Survey on Multilingual Mental Disorders Detection
Ana-Maria Bucur1,2,3, Marcos Zampieri4, Tharindu Ranasinghe5, Fabio Crestani1 1Università della Svizzera italiana, Switzerland 2Interdisciplinary School of Doctoral Studies, University of Bucharest, Romania 3PRHLT Research Center, Universitat Politècnica de València, Spain 4George Mason University, USA 5Lancaster University, UK ana-maria.bucur-cosma@usi.ch
@@ 1 Introduction
Nearly half of the global population risks developing at least one mental disorder by age 75 McGrath et al. (2023). Many individuals avoid seeking help due to stigma, which varies across cultures and is shaped by cultural norms, religious beliefs, and social attitudes Ahad et al. (2023). Due to several challenges, including the stigma surrounding mental health, limited access in certain areas, economic factors, and a high global demand for services, the World Health Organization advocates for improved delivery of mental health services, including digital technologies to deliver remote care. [^1] https://www.who.int/news/item/17-06-2022-who-highlights-urgent-need-to-transform-mental-health-and-mental-health-care Furthermore, integrating remote screening tools and culturally adapted digital interventions is crucial Bond et al. (2023), as remote screening can detect language patterns linked to mental disorders from short essays Rude et al. (2004), text messages Nobles et al. (2018), or social media Eichstaedt et al. (2018). However, developing effective screening tools for mental disorders heavily relies on the availability of data. Therefore, in this work, we present the datasets that can be used to train models for screening in languages other than English.
@@
The first well-known study on the detection of mental disorders using social media was conducted by De Choudhury et al. (2013). Subsequent research has shown that the language used on Facebook can predict future depression diagnoses found in medical records, indicating that social media data could serve as a valuable complement to depression screening Eichstaedt et al. (2018). Methods used for social media screening focus mainly on English data Skaik and Inkpen (2020); Harrigian et al. (2021), and workshops and shared tasks addressing NLP applications to mental health, such as eRisk Parapar et al. (2024); Crestani et al. (2022), CLPsych Chim et al. (2024) and LT-EDI Kayalvizhi et al. (2023) also primarily use English data.
@@
Current NLP models face major limitations in mental disorders detection in languages other than English. Studies show that cultural differences shape how mental disorders are expressed in language De Choudhury et al. (2017); Aguirre and Dredze (2021); Rai et al. (2024), which means that markers predictive in English often do not generalize well across different cultures Aguirre et al. (2021); Abdelkadir et al. (2024). Even one of the best predictors of depression in language, the use of the first person pronoun “I” Rude et al. (2004), for example, has different degrees of association with the severity of depression across different demographic groups Rai et al. (2024). This suggests that markers of mental disorders in language are not universal. In addition, self-disclosure rates vary across cultures; collectivist cultures tend to exhibit lower self-disclosure rates than individualist cultures in online settings Tokunaga (2009). Furthermore, non-native English speakers tend to use their native language for more intimate self-disclosures on social media, with higher rates of negative disclosure compared to posts in English Tang et al. (2011). This could have substantial implications for English-based social media screening tools, as they may overlook signals of mental disorders that are present in posts written in languages other than English.
@@
Recently, there have been efforts to develop detection models that focus on languages other than English, such as Portuguese Santos et al. (2024), German Zanwar et al. (2023), Arabic Almouzini et al. (2019), and Chinese Zhu et al. (2024). There have also been shared tasks specifically designed to address these issues, such as MentalRiskES Mármol-Romero et al. (2025), which focuses on the early detection of depression, suicide, and eating disorders in Spanish. To further contribute to these important efforts, we present the first survey on the detection of mental disorders from social media data beyond English. This survey aims to promote the development of multilingual NLP models that account for cross-cultural and cross-linguistic differences in online language.
@@
    We provide a comprehensive list of multilingual mental health datasets that capture linguistic diversity and can be used for developing multilingual NLP models. [^2] We make the list publicly available, and we will continuously update it: https://github.com/bucuram/multilingual-mental-health-datasets-nlp
@@
    We discuss cross-cultural and cross-language differences in the manifestations of mental disorders in social media.
@@
    We identify and describe several research gaps and future directions in the detection of mental disorders using online data beyond English.
@@ 2 Related Surveys
In this section, we analyze related surveys on the analysis of mental disorders from social media data. Calvo et al. (2017) is considered one of the first comprehensive surveys, presenting the datasets and NLP techniques used for mental health status detection and intervention. The survey explores research on various mental health conditions and states, including depression, mood disorders, psychological distress, and suicidal ideation, specifically in non-clinical texts such as user-generated content from social media and online forums. Similarly, recent surveys Skaik and Inkpen (2020); Harrigian et al. (2021); Ríssola et al. (2021); Zhang et al. (2022); Garg (2023); Bucur et al. (2025a, b) present the datasets, features, and models used to detect mental disorders from online content, with a primary focus on English language data. Dhelim et al. (2023) and Bucur et al. (2025b) focus on studies that were published during the COVID-19 pandemic that address mental well-being, loneliness, anxiety, stress, PTSD, depression, suicide, and other mental disorders. In addition to these surveys, Chancellor and De Choudhury (2020) provides a critical review of the study design and methods used to predict mental health status, along with recommendations to improve research in this field.
Our paper fills an important gap in the literature by offering the first comprehensive survey of research on detecting mental disorders in languages other than English. The most related survey to ours is the one conducted by Garg (2024), which focuses solely on low-resource languages and only discusses datasets in Thai, Bengali, Hindi, Japanese, and Korean. In contrast, our survey has a broader scope, covering research on a variety of languages, regardless of their resource availability. Our work complements previous surveys by providing an important data-centric foundation for researchers seeking to deploy and validate NLP methods in non-English contexts.
@@ 3 Mental Disorders Detection Tasks Overview
In this section, we discuss common tasks related to predicting mental disorders, emphasizing research available in non-English languages where applicable. The prediction of mental disorders via social media typically involves a supervised binary classification task, in which online posts are used to determine the presence or absence of disorders (Figure 1). This can be performed at the post level for predicting suicidal ideation Huang et al. (2019) and depression Uddin et al. (2019); Bucur et al. (2023), or at the user level for conditions like depression Hiraga (2017), anxiety Zarate et al. (2023), and bipolar disorder Sekulić et al. (2018). User-level prediction can serve as an early risk prediction task, facilitating timely assessments Losada and Crestani (2016); Bucur et al. (2021); Mármol-Romero et al. (2023). Severity prediction assesses the intensity of conditions such as depression Naseem et al. (2022); Kabir et al. (2023); Sampath and Durairaj (2022) or suicide risk Benjachairat et al. (2024). Longitudinal analysis of social media helps detect moments of change in mental health Tsakalidis et al. (2022). Other tasks focus on symptom prediction Liu et al. (2023); Yadav et al. (2020) and highlighting evidence for mental health problems Chim et al. (2024); Varadarajan et al. (2024). Mental health indicators from the social media timeline of an individual can be used to fill in validated questionnaires, with the goal of estimating symptoms of mental disorders that are usually assessed through survey-based methods such as BDI-II [^3] https://naviauxlab.ucsd.edu/wp-content/uploads/2020/09/BDI21.pdf for depression assessment Parapar et al. (2021) or EDE-Q [^4] https://www.corc.uk.net/media/1273/ede-q_quesionnaire.pdf for eating disorders Parapar et al. (2024). Lastly, mental health monitoring uses aggregated detection results to estimate the prevalence of disorders, as demonstrated during the COVID-19 pandemic Cohrdes et al. (2021).
@@
Shared tasks have encouraged interdisciplinary collaborations between psychologists and computer scientists. English-based shared tasks, such as eRisk Parapar et al. (2024), CLPsych Coppersmith et al. (2015), and LT-EDI Kayalvizhi et al. (2023) provided valuable benchmark datasets that the research community continues to use, even beyond the official competitions. MentalRiskES [^5] https://sites.google.com/view/mentalriskes2025 Mármol-Romero et al. (2025) is the only shared task focused on detecting mental disorders in languages other than English. MentalRiskES includes tasks such as the detection of depression, anxiety, eating disorders, and suicidal risk in the Spanish language Mármol-Romero et al. (2023).
4 Methodology
Figure 2: PRISMA flow diagram for our survey.
@@ 4 Methodology
Sinhala Low Rathnayake and Arachchige (2021), EmoMent Atapattu et al. (2022), Herath and Wijayasiriwardhane (2024)
The classification of Sinhala is based on the work of Ranathunga and de Silva (2022), who identified the miscategorization of Sinhala in Joshi et al. (2020)’s classification.
Table 1: Datasets for detecting mental disorders in languages other than English included in the current survey. The availability of resources is based on Joshi et al. (2020).
Our primary objective is to catalog the available non-English resources derived from social media for the detection of mental disorders. We conducted a systematic search on major publication databases, including ACL Anthology, ACM Digital Library, IEEE Xplore, Springer Nature Link, ScienceDirect, and Google Scholar. Initially, 405 studies were identified through database searches. After screening the abstracts, 215 papers were excluded because they did not mention the language of the data or mention that the data was in English. Thus, following a review of the main body of the papers, the number of eligible studies was narrowed down to 108, which represents the final count of papers included in this survey. Papers that did not present new data collections in languages other than English were excluded during the screening process. The PRISMA flow diagram is presented in Figure 2. The search terms used in our survey are provided in the Appendix A.
@@ 5 Multilingual Datasets
In this section, we outline the datasets included in the current survey. The languages most frequently represented in these data collections are three high-resource languages: Chinese, Arabic, and Spanish. Although approximately half of the datasets were published in unranked venues, which diminishes their visibility, the other half were published in high-ranking journals and conferences (Figure 4 in Appendix A.
@@ 5.1 Data Sources
Most of the datasets in English are sourced from Twitter [^6] All the datasets were collected before Twitter changed its name to X, so we refer to it as ‘Twitter’ in this paper. and Reddit Harrigian et al. (2021). Most non-English datasets in this section were also primarily collected from Twitter. However, Reddit was not as widely used for these data collections in non-English contexts. The data presented in this survey come from various populations and regions, and some of the sources are platforms that are exclusive to specific countries, such as Sina Weibo [^7] https://weibo.com used in China, VKontakte [^8] https://vk.com/ used in Russia, Pantip [^9] https://pantip.com/ in Thailand, or Everytime [^10] https://everytime.kr/ in Korea. Complete information about the sources of the data is presented in Table LABEL:tab:datasets in Appendix A.
The way individuals use social media platforms to share information about themselves varies not only by platform but also by culture. Twitter provides community and safety, helping raise awareness and combat stigma around mental health Berry et al. (2017). In contrast, Reddit allows for greater anonymity with “throwaway” accounts, encouraging users to openly share their experiences in detailed posts on specific subreddits De Choudhury and De (2014). This longer format supports post-level mental health analysis Chowdhury et al. (2024), while Twitter’s shorter posts favor user-level insights, requiring longitudinal data to identify language patterns Tumaliuan et al. (2024). Moreover, even when individuals come from the same cultural background and speak the same language, the topics they choose to discuss are heavily influenced by the platform they use. For example, on Sina Weibo, users primarily talk about popular culture, while on Twitter, conversations in Chinese revolve more around politics Zhang and Gonçalves (2016). These differences, potentially shaped by platform governance and censorship, have direct implications for mental health data collection, as they influence the visibility, framing, and prevalence of mental health-related discourse across different platforms.
@@ 5.2 Languages
Table 1 presents all the datasets included in this survey. To classify the availability of resource types, we used the framework proposed by Joshi et al. (2020), which categorizes resources into six classes (from Rare to High) based on the availability of both labeled and unlabeled NLP data. The languages most frequently represented in the data collections are high-resource languages: Chinese appears in 25 data collections, Arabic is found in 11 datasets, and Spanish is included in 10 datasets. Even if most of the languages covered in the data are from high-, mid to high- and mid-resourced languages, we also have some languages with fewer resources, such as Cantonese, Norwegian and Sinhala. Most of the languages used in the data collections belong to some of the largest language families by number of speakers, specifically the Indo-European, Sino-Tibetan, and Afro-Asiatic language families. When analyzing the language resources based on speaker population and vitality according to Ethnologue Eberhard et al. (2023), we find that all the languages listed in Table 1 are categorized as large-institutional. This classification indicates that these languages are supported and maintained by institutions beyond their respective communities and have a large number of speakers. The categorization of all listed languages as large-institutional by Ethnologue underscores a robust support system that enhances their viability in NLP applications. However, it also highlights the urgent need to address disparities in resource availability for less-represented languages, promoting inclusivity in NLP research.